The coefficient of kinetic friction for a 22kg bobsled on a track is 0.1, what force is required to push down a 6 degree incline an achieve a speed of 60 km/h at the end of 75m
1
Expert's answer
2020-11-25T10:50:41-0500

From Newton's second law:


"F + mg\sinθ - μmg\cosθ = ma\\F =\frac{mv^2}{2s}- mg\sinθ + μmg\cosθ \\F =\frac{(22)(60)^2}{2(75)3.6^2}- (22)(9.8)\sin{6} \\ + (0.1)(22)(9.8)\cos{6} =39.6\ N"